The following Table-3 depicts the market for rice (millions of kilos). Use this table to

answer following questions.

3

Table-3

 Price ($) of rice

 2.00

 4.00

 6.00

 8.00

 10.00

 Quantity demanded

 90

 80

 70

 60

 50

 Quantity supplied

  30

  50

  70

  90

  110

 (a) Draw the demand and the supply curves for rice.

(b) What is the present equilibrium price and quantity exchanged in this market?

(c) Suppose that government introduces a price floor of $8 kilo. How much in total will

rice buyers now be paying?

(d) As a result of the price floor, what will be the total amount of surplus? What will be

the dollar amount of this surplus?

(e) Who will be responsible for buying this surplus?


1
Expert's answer
2021-10-07T10:02:48-0400

(a)



(b)

The present equilibrium price =$6

The present equilibrium quantity= 70.


(c)

Price floor=$8

Equilibrium quantity demanded=7

"\therefore" in total, buyers will be paying

"8\times70 = 560"


(d)

Surplus will be given by the area of the triangle formed below the equilibrium level and above the supply curve:

"= \frac{1}{2}\times(70-30)\times(6-2)"

"=80."

Dollar amount:

"=8\times80"

=$640.


(e)

The government will be responsible for buying the surplus.